OpenStudy (kayluh123):

When the air in a hot-air balloon is heated to 100 degrees Fahrenheit, each cubic foot (ft3) of air can lift about 7 grams of weight.To lift 1000 pounds, you must have ________ ft3 of air. Round answer to the nearest cubic foo

OpenStudy (anonymous):

First convert pounds to grams. 1000lbs = 453 592.37g Set up a proportion. 7g/1ft3 = 453 592.37g/x {divide by 453 592.37g} 1.54323584 × 10-5/ft3=1/x {inverse both sides} 64 798.91ft3 = x Round and delable = 64,799